# Guiding In-Context Learning of LLMs through Quality Estimation for Machine Translation

### v1

@@ 1 Introduction
Naturally, the most effective ICEs for a given source text are the ones that would impact the resulting translation quality, which would ultimately depend on translation references or human judgment. In MT, quality estimation (QE) has become a standard approach for evaluating an MT system’s output without relying on reference translations Blain et al., (2023). Recently, Lee, (2020), Ye and Li, (2023), and Sharami et al., 2023a showed the effectiveness of domain-specific QE when it comes to domain-specific MT (in contrast to the ineffectiveness of generic QE). Building on this and to address the aforementioned challenges, our work proposes to leverage domain-specific QE to assist in the selection of ICEs, with the goal of determining the suboptimal number and combination of ICEs to maximize MT quality, all without reference translations. As QE would assess the impact of different ICE combinations and sequences, we hypothesize that this integration has the potential to not only improve translation performance but also reduce processing time, as QE could result in smaller sets of ICEs, which would reduce the inference times (Petrov et al.,, 2023). This is particularly crucial considering the limited number of ICEs that can be fed into LLMs (Agrawal et al.,, 2023). Therefore, our study aims to investigate the feasibility of selecting ICEs on a per-source basis. Specifically, we aim to answer the following research question (RQ): How effective are domain-specific QE models in determining ICEs for translation tasks in an LLM?
@@ 3.2 Quality Estimation
Following (Ranasinghe et al.,, 2020; Lee,, 2020; Sharami et al., 2023a, ), we develop a domain-specific QE model. First, we trained a QE model using out-of-domain (OOD) data (as detailed in Section 3.2.1) to ensure generalizability; and second, we fine-tuned the model using the training set described in Section 3.4 to provide domain-specific QE model and address domain mismatch, which is critical (Koehn and Knowles,, 2017).
@@ 3.2.1 QE data
Using SacreBLEU, we calculated the BLEU score, comparing the translated text with its corresponding post-edited text. This approach, which has been demonstrated to be effective in QE (Negri et al.,, 2018; Lee,, 2020; Sharami et al., 2023a, ), enabled us to use the source and (machine-) translated text as input and the BLEU score as the target value for the QE model. For building domain-specific QE, we utilized the training set detailed in Section 3.4 and applied the aforementioned approach to synthetically generate BLEU scores for the entire dataset.
@@ 3.7 Computational Costs
The script for running our experiments is publicly available at the GitHub repository.
@@ 4.1 Time to Prediction (TTP)
It is also important to note the time required to train the QE models used in the prediction process. As provided in Appendix 3, the training time for the generic QE model is +/+/-+ / - 5 hours and 55 minutes, while the specific QE model takes about 6 hours and 54 minutes. Although these training times are significant, it is crucial to recognize that QE models, similar to MT models, can be reused for the same language pair and domain, thereby amortizing the initial training cost over multiple predictions.
@@ 6 Related Work
As shown by Vilar et al., (2023), the quality of translation is directly proportionate to the quality of ICEs, where quality refers to ICEs being relevant, clear, accurate, and domain-specific. However, considering all pairs of ICEs during processing is computationally demanding (Alves et al.,, 2023). Hence, it is crucial to selectively choose ICEs that can enhance MT quality. Goyal et al., (2022) conducted a study where ICEs were randomly selected. Despite finding that this random selection of ICEs resulted in good translation performance, the neglect of their order, which was identified as important (Liu et al.,, 2022; Lu et al.,, 2022), was a drawback in this approach. To address the order issue, methodologies such as (Agrawal et al.,, 2023) introduced a re-ranking technique (R-BM25). However, their methodology relies solely on n-grams to order examples, which can enhance fluency but may overlook contextual factors. In our proposed ICL approach, we investigated the unigram order of initial ICEs provided by the BM25 algorithm. We leave the in-depth analysis of ICE order for future work.
@@
QE models offer a quick solution to the assessment of the overall usefulness of translated text. These models do not rely on reference translations, thereby reducing the human effort required for quality evaluation (Tamchyna,, 2021; Murgolo et al.,, 2022; Zerva et al.,, 2022; Blain et al.,, 2023). Similar to MT models, previous studies highlight the importance of domain-specific QE for accurately estimating translation quality across diverse domains (Lee,, 2020; Sharami et al., 2023b, ). This is why, in our work, we employed a domain-specific QE model instead of a generic one to enhance the selection of ICEs.
@@ 7 Conclusion
We evaluated our approach across different modes and early stopping patience values on the German-to-English IT dataset. Our experiments consistently demonstrated the superior performance of our methodology, surpassing all prior works across both BLEU and COMET metrics. Our method consistently improves BLEU scores, although this comes at the cost of increased computation time. We also investigated the impact of ordering the ICEs based on their unigram overlap with the source text and found it to be not statistically significant. Furthermore, our experiments highlighted the value of ICL compared to fine-tuning a pre-trained large model, namely mBART-50. We also highlighted that our method leads to less carbon emissions while achieving better translation performance.
In the future, we would like to conduct further research on the impact of our proposed methodology across different datasets and LLMs. Additionally, we aim to explore alternative metrics beyond BLEU to tailor the selection process, as well as additional features such as bigram, type/token ratio, and length when ordering examples prior to their input into LLMs.

### v2

@@ 1 Introduction
Naturally, the most effective ICEs for a given source text are the ones that would impact the resulting translation quality, which would ultimately depend on translation references or human judgment. In MT, quality estimation (QE) has become a standard approach for evaluating an MT system’s output without relying on reference translations Blain et al., (2023). Recently, Lee, (2020), Ye and Li, (2023), and Sharami et al., (2023) showed the effectiveness of domain-specific QE when it comes to domain-specific MT (in contrast to the ineffectiveness of generic QE). Building on this and to address the aforementioned challenges, our work proposes to leverage domain-specific QE to assist in the selection of ICEs, with the goal of determining the suboptimal number and combination of ICEs to maximize MT quality, all without reference translations. As QE would assess the impact of different ICE combinations and sequences, we hypothesize that this integration has the potential to not only improve translation performance but also reduce processing time, as QE could result in smaller sets of ICEs, which would reduce the inference times (Petrov et al.,, 2023). This is particularly crucial considering the limited number of ICEs that can be fed into LLMs (Agrawal et al.,, 2023). Therefore, our study aims to investigate the feasibility of selecting ICEs on a per-source basis. Specifically, we aim to answer the following research question (RQ): How effective are domain-specific QE models in determining ICEs for translation tasks in an LLM?
@@ 3.2 Quality Estimation
Following (Ranasinghe et al.,, 2020; Lee,, 2020; Sharami et al.,, 2023), we develop a domain-specific QE model. First, we trained a QE model using out-of-domain (OOD) data (as detailed in Section 3.2.1) to ensure generalizability; and second, we fine-tuned the model using the training set described in Section 3.4 to provide domain-specific QE model and address domain mismatch, which is critical (Koehn and Knowles,, 2017).
@@ 3.2.1 QE data
Using SacreBLEU, we calculated the BLEU score, comparing the translated text with its corresponding post-edited text. This approach, which has been demonstrated to be effective in QE (Negri et al.,, 2018; Lee,, 2020; Sharami et al.,, 2023), enabled us to use the source and (machine-) translated text as input and the BLEU score as the target value for the QE model. For building domain-specific QE, we utilized the training set detailed in Section 3.4 and applied the aforementioned approach to synthetically generate BLEU scores for the entire dataset.
@@ 3.7 Computational Costs
The script for running our experiments is publicly available at anynomous.com.
@@ 4.1 Time to Prediction (TTP)
It is also important to note the time required to train the QE models used in the prediction process. As provided in Appendix 3, the training time for the generic QE model is +/+/-+ / - 5 hours and 55 minutes, while the specific QE model takes about +/+/-+ / - 6 hours and 54 minutes. Although these training times are significant, it is crucial to recognize that QE models, similar to MT models, can be reused for the same language pair and domain, thereby amortizing the initial training cost over multiple predictions.
@@ 6 Related Work
As shown by Vilar et al., (2023), the quality of translation is directly proportionate to the quality of ICEs, where quality refers to ICEs being relevant, clear, accurate, and domain-specific. However, considering all ICEs during processing is computationally demanding (Alves et al.,, 2023). Hence, it is crucial to selectively choose ICEs that can enhance MT quality. Goyal et al., (2022) conducted a study where ICEs were randomly selected. Despite finding that this random selection of ICEs resulted in good translation performance, the neglect of their order, which was identified as important (Liu et al.,, 2022; Lu et al.,, 2022), was a drawback in this approach. To address this, methodologies such as (Agrawal et al.,, 2023) introduced a re-ranking technique (R-BM25). However, their methodology relies solely on n-grams to order examples, which can enhance fluency but may overlook contextual factors. In our approach, we investigated the unigram order of initial ICEs provided by the BM25 algorithm. We leave the in-depth analysis of ICE order for future work. Additionally, Kumar et al., (2023) highlighted the advantages of using multiple features in ICE selection to improve translation quality, while our QE-based approach simplifies ICE selection without needing to generate additional features, ensuring efficiency.
@@
QE models offer a quick solution to the assessment of the overall usefulness of translated text. These models do not rely on reference translations, thereby reducing the human effort required for quality evaluation (Tamchyna,, 2021; Murgolo et al.,, 2022; Zerva et al.,, 2022; Blain et al.,, 2023). Similar to MT models, previous studies highlight the importance of domain-specific QE for accurately estimating translation quality across diverse domains (Lee,, 2020; Sharami et al.,, 2023). This is why, in our work, we employed a domain-specific QE model instead of a generic one to enhance the selection of ICEs.
@@ 7 Conclusion
We evaluated our approach across different modes and early stopping patience values on the German-to-English IT dataset. Our experiments consistently showed the superior performance of our methodology, surpassing all prior works across both BLEU and COMET metrics. Our method consistently improves BLEU scores, although this comes at the cost of increased computation time. We also investigated the impact of ordering the ICEs based on their unigram overlap with the source text and found it to be not statistically significant. Furthermore, our experiments highlighted the value of ICL compared to fine-tuning a pre-trained large model, namely mBART-50. We also highlighted that our method leads to less carbon emissions while achieving better translation performance.
In the future, we would like to conduct further research on the impact of our proposed methodology across different language pairs, domains and LLMs. Also, we aim to explore alternative metrics beyond BLEU to tailor the selection process, as well as additional features such as bigram, type/token ratio, and length when ordering examples prior to their input into LLMs.
8 Acknowledgments
We would like to thank the anonymous reviewers for their helpful comments.
Appendices
Figure 2: Overview illustration showing an iteration of our proposed methodology.
Metric Generic Model Specific Model
Training Time (hh:mm) 05:55 06:54
CO2 Emissions (kg) 1.41 1.46
Electricity Consumption (kWh) 3.63 3.76
Table 3: Training Time, CO2 Emissions, and Electricity Consumption for QE Models.
ICEs:
Die Sockets, die im except Array aufgelistet sind, werden auf Ausnahmen überwacht. = The sockets listed in the except array will be watched for exceptions. $</s>$ Geben Sie den Namen der Variablen ein, deren Wert überwacht werden soll. = Enter the name of the variable whose value is to be monitored. $</s>$ Nur erlaubt bei Sockets für lokale Displays und den globalen Socket. = Permitted only on sockets of local displays and the global socket. $</s>$ Legt fest, ob Scandaten-Information, die in den MPEG2-Videoströmen enthalten sind, aktualisiert werden sollen. = This controls whether to update the scan data information contained in the MPEG-2 video streams. $</s>$ Die Sockets, die im write Array aufgelistet sind, werden daraufhin überwacht, ob ein Schreibvorgang den Socket blockiert. =
Translation:
The sockets listed in the write array will be watched for whether a write operation blocks the socket.
Reference Label:
The sockets listed in the write array will be watched to see if a write will not block.
QE: 67.59, BLEU score (using reference label): 52.89
Table 4: An example of selected ICEs for a source text, its corresponding translation, reference label, and QE estimation compared to the BLEU score computed based on the reference label.
1:function Search(…)
2:     $\text{temp}\leftarrow[(``",0.0,``")]$
3:     $\text{prompt}\leftarrow``"$
4:     $\text{itr}\leftarrow 0$
[…]
9:         if available_prompts is not empty then
[…]
15:                  return temp
16:              end if
17:              $\text{output}\leftarrow\textsc{GenerateOutput}(...)$
[…]
23:                  return temp
24:              end if
25:              $\text{temp}\leftarrow\textsc{SortTemp}(...)$
[…]
28:              else
[…]
30:              end if
[…]
32:         end if
33:         $\text{itr}\leftarrow\text{itr}+1$
34:     end while
35:     return temp
36:end function
Algorithm 1 Pseudocode outlining the proposed Search Algorithm. Each phase of the methodology is annotated alongside the relevant code. Function arguments are omitted for simplicity. The first element of the returning list (temp) includes the selected prompt, its associated QE score, and the translated text.